OpenStudy (appleduardo):

how to solve (x^2 + y^2)dx + xy dy= 0 ??

OpenStudy (appleduardo):

im stuck in : dy/dx = x/y + y/x + 1/y

OpenStudy (anonymous):

Wolframalpha if no one answers

OpenStudy (appleduardo):

\[\frac{ dy }{ dx } = \frac{ x }{ y} + \frac{ y }{ x } + \frac{ 1 }{ y }\]

OpenStudy (appleduardo):

what to do now?

OpenStudy (anonymous):

? your goal is to get y`, and you've done it